OpenStudy (anonymous):

Which of the following is equivalent to the proportion 2/3 = 3/y? A. 3y = (2)(3) B. 2y = (2)(3) C. 2y = (3)(3) D. 3y = (3)(3)

OpenStudy (anonymous):

By cross multiplication, 2y = (3)(3) so C

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@s6cppr Do you mind helping me with a few more similar problems?

OpenStudy (anonymous):

If it seems vague then: 2/3 = 3/y multiply 3 both sides to cancel 3 on the left so 3*(2/3) = 2 2= 3*(3/y) 2 = 9/y multiply by y again to simplify right hand side so 2y = 9 or (3)(3)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which of the following is equivalent to the proportion 3/4 =x/12? a. 4/3 =x/12 b. 3/4 = 12/x c. 4/3=x/12 d. 4/12=x/3

OpenStudy (anonymous):

cross multiply to get 36 = 4x which is 9. Apparently, 3/4 x (3/3) = 9/12 so: B as 3/4 x (4/4) = 12/16 3/3 = 4/4 = 1

OpenStudy (anonymous):

@s6cppr Alright, got it. Thanks again for all your help :)

OpenStudy (anonymous):

np